OpenStudy (anonymous):

Celia is staring at the clock waiting for school to end so that she can go to track practice. She notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 1:25 to 1:50? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? You must show all of your work.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Okay. Part 1: Degrees per minute = 360 degrees/60 minutes = 6 degrees/min. From 1:25 to 1:50, 25 minutes has passed, so the minute hand has moved 25*6 = 150 degrees or 150*pi/180 = 5pi/6 radians. Part 2: Total distance the tip of minute hand has traveled in this time = 2pi*(4 in) * (5pi/6)/(2pi) = 10pi/3 inches.

OpenStudy (anonymous):

Hope this helped! Have a great day! @Jaylyn.Ki

OpenStudy (anonymous):

LOVE YOU SO MUCH!